OpenStudy (anonymous):

The equation shows the relationship between x and y: y = −2x + 11 What is the slope of the equation? −11 −2 9 11

OpenStudy (anonymous):

@dan815 @Daniellelovee

OpenStudy (anonymous):

Hey!

OpenStudy (anonymous):

hii

OpenStudy (anonymous):

y = -2x + 11 is in the form of y = mx + b y = mx + b is called the slope intercept form! In y= mx + b, m is the slope and b is the y-intercept!

OpenStudy (anonymous):

whoa lol give me a sec to try and understand lol :D

OpenStudy (anonymous):

Sure! \(\bf y = \color{red}{m}x + \color{green}{b}\) \(\bf \color{red}{m} = slope\) \(\bf\color{green}{b} = y-intercept\)

OpenStudy (anonymous):

oh so the slope is -2?

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

thanks!!! could you help with some more please? ill repost :)